### v1

# Adversarial Robustness in Parameter-Space Classifiers

Abstract
Implicit Neural Representations (INRs) have been recently garnering increasing interest in various research fields, mainly due to their ability to represent large, complex data in a compact and continuous manner. Past work further showed that numerous popular downstream tasks can be performed directly in the INR parameter-space. Doing so can substantially reduce the computational resources required to process the represented data in their native domain. A major difficulty in using modern machine-learning approaches, is their high susceptibility to adversarial attacks, which have been shown to greatly limit the reliability and applicability of such methods in a wide range of settings. In this work, we show that parameter-space models trained for classification are inherently robust to adversarial attacks – without the need of any robust training. To support our claims, we develop a novel suite of adversarial attacks targeting parameter-space classifiers, and furthermore analyze practical considerations of attacking parameter-space classifiers. Code for reproducing all experiments and implementation of all proposed methods will be released upon publication.
1 Introduction
Implicit Neural Representations (INRs) are representations of an arbitrary signal as a neural network predicting signal values under some conditioning. These representations have recently been raising increased interest due to their ability to efficiently and compactly encode high-dimensional data while maintaining high input fidelity, and have further shown to be capable of learning additional intrinsic knowledge latent within the original signal-domain representation (Sitzmann et al., 2020; Wang et al., 2021).
Figure 1: t-SNE projection of modulation vectors fitted for clean and adversarialy-perturbed data - for parameter-space and signal-space classifiers.
While INRs are inherently designed to fit one specific signal (rather than generalizing across an entire dataset), several recent works have proposed ways to use these representations to achieve a variety of downstream tasks by using some neural meta-network operating directly over the parameter-space of the INR of each signal within the dataset ((Chen & Wang, 2022; Dupont et al., 2022; Lee et al., 2021; Navon et al., 2023b)).
In these settings, each data sample is received in the native signal space, converted to its compressed INR representation by fitting a neural model, and finally the meta-network performs the designated downstream task over the implicit representation encoded by the model’s parameters. The main advantage of such approaches is that they alleviate challenges in tasks involving high-dimensional data (videos, point-clouds, spectrograms), often requiring the incorporation of large, complex and problem-specific neural architectures for efficient data processing, and also potentially demand large computational resources for efficient optimization.
Adversarial attacks are known to be a major concern for the use of neural networks across a variety of tasks. In this work, we explore – to the best of our knowledge, for the first time – the adversarial robustness of parameter-space classifiers. We start by framing the problem: any attack on a parameter-space classifier must be simultaneously optimized to reduce classification performance while also satisfy some constraint incurring the original, signal domain data fidelity. Since no prior studies have addressed this setting, we first formulate a novel set of possible white-box attacks over several data modalities. We then proceed to show that discussed parameter-space classification models are inherently robust to adversarial attacks in comparison to the signal-domain classification problem counterpart, as exemplified using our established attack suites.
Our hypothesis of inherent robustness can be motivated by the following: Firstly, adversarial attacks commonly utilize the highly-nonlinear nature of modern deep classifiers to create perturbations with specific patterns that, while bounded, will be heavily amplified by the classifier to create significant data manipulations in deeper layers ((Ingle & Kulkarni, 2021; Romano et al., 2020)). In this setting of parameter-space classification, INR optimization is performed per-sample, and is embedded onto the classifier’s inference process. This constrains the attacker to optimize perturbations in the signal-domain rather than directly manipulating the classifier’s input. This optimization is disjoint from the downstream objective, and is dedicated to expressing global data features as a whole, rather than entangle specific entries with specific local signal features (Sitzmann et al., 2020). Therefore, this optimization process tends to obfuscate adversarial patterns introduced in signal domain. Secondly, as we later show in Section 4.4, performing adversarial attacks over parameter-space classifiers poses significant computational difficulties for a white-box attacker, as attacking the INR given a perturbed signal requires backpropagation through an internal optimization loop.
In summary, this paper makes the following contributions:
  1. 1.
    We supply evidence of inherent robustness of parameter-space classifiers to adversarial attacks.
  2. 2.
    We propose several novel types of adversarial white-box attacks for parameter-space classifiers.
  3. 3.
    To demonstrate our claims on 3D data, we also formulate a novel adversarial attack for voxel-grid data representations, and demonstrate its performance in both parameter-space and signal-space classification settings.
2 Related Work
The analysis of trained neural network parameters had been raising increasing interest in recent years, demonstrating applicability in a wide, and at times surprising, range of machine learning tasks. (Eilertsen et al., 2020) was one of the first works considering this idea, performing trained network parameter analysis to identify the optimization procedure hyperparameters. (Unterthiner et al., 2020) showed that a trained CNN classifier model parameters can be used to directly attain a good assessment of classification accuracy, without running the full model’s forward pass. (Dupont et al., 2022) have been the first to demonstrate that neural network parameters, specifically of the SIREN (Sitzmann et al., 2020) models (one of the first INRs to be proposed), could be used as the direct and only input for the training of a meta-learner model for classification, generative, data-imputation and novel-view synthesis tasks. (Peebles et al., 2022) employ generative modeling of deep-weight space to improve parameter update steps in model training.
3 Method
3.1 Implicit neural representations
Let $\mathcal{X}$ denote a class of signals of the form $x:\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ (for clarity, we consider the domain and the co-domain to be finite Euclidean spaces; a more general formulation is straightforward). Given a specific signal $x$, the fitting of an implicit neural representation consists of fitting a neural model $F_{\theta}$ parameterized by $\theta\in\mathcal{Y}$ so to approximate $x$ in the sense of some norm, (Chen & Wang, 2022; Dupont et al., 2022)
$\theta^{*}(x)=\operatorname*{arg\,min}_{\theta\in\mathcal{Y}}\quad$ $\|F_{\theta}-x\|_{\mathcal{X}}.$ (1)
In what follows, we denote the representation obtained by solving the minimization problem in equation (1) as a map $R:\mathcal{X}\to\mathcal{Y}$, so that $R:x\mapsto\theta^{*}(x)$.
3.2 Parameter-Space Classifiers
In contrast to conventional classifiers operating directly on the signal space $\mathcal{X}$, parameter-space classifiers operate on the signal INRs, $R(\mathcal{X})\subset\mathcal{Y}$. We denote a parameter-space classifier as $M_{\psi}$, parameterized by $\psi$; the model receives the representation parameters $\theta=R(x)$ as the input and outputs the predicted class label $\hat{c}=M_{\psi}(\theta)$ as the output.
Even for small neural representation models, the parameter space $\mathcal{Y}$ is often high-dimensional, increasing difficulties related to computational resources and overfitting when training the classification models on it (Navon et al., 2023a, b). Functas (Dupont et al., 2022) cope with this issue of increased parameter-space dimensionality by partitioning the parameter space into weights and biases, $\mathcal{Y}=\mathcal{Y}_{\mathrm{w}}\cup\mathcal{Y}_{\mathrm{b}}$, and sharing a single vector of weights $\theta_{\mathrm{w}}\in\mathcal{Y}_{\mathrm{w}}$ across a collection of represented signals, while allowing to fit individual modulation vectors $v(x)\in\mathbb{R}^{d}$ of a fixed dimension with a learned shared linear decoder $W$ such that each signal has an individual vector of biases, $\theta_{\mathrm{b}}(x)=Wv(x)$. This technique was successfully demonstrated with the SIREN (Sitzmann et al., 2020) models.
In what follows, we continue denoting the representation of $x$ as $\theta$, while taking into account that the actual INR $F$ may have additional parameters shared across multiple signals.
3.3 Parameter-Space Adversarial Attacks
Given a pre-trained parameter-space classification model $M_{\psi}$, our objective is to optimize an adversarial perturbation $\delta$ in signal domain under some $L_{q}(\mathcal{X})$-norm bound,
$\max_{\delta}\quad$ $d(M_{\psi}(R(x+\delta)),M_{\psi}(R(x)))\,\,\,\text{s.t.}\,\,\,\|\delta\|_{L_{q}(\mathcal{X})}\leq\epsilon,$ (2)
where $d$ denotes a distance in the class label domain.
Note that while the signal is attacked in its native domain, the classifier $M$ accepts its input in the representation domain. The two domains are related by the map $R$ – a function whose pointwise value requires the solution of an optimization problem via a numerical algorithm. While backpropagating through this optimization process is possible (e.g. via automatic differentiation or implicit differentiation (Navon et al., 2020)), as we later show in Section 4, doing so dramatically increases computational resources required for perturbation optimization, thus posing additional practical challenges for a potential attacker, in addition to the inherent difficulty of optimization such an attack that stems from the inference pipeline itself, as motivated in section 1.
In the following sections, we propose five different adversarial attacks, addressing these challenges that are absent in the traditional signal-domain adversarial attack settings.
3.3.1 Full PGD
Figure 2: Parameter-space classifier adversarial attack pipeline - for a single PGD iteration. Orange blocks are optimized, blue blocks remain frozen.
A straightforward approach to solving equation (2) is by performing projected gradient descent (PGD) and backpropagating through both the classification model $M$ and the modulation optimization steps collectively denoted $R(x)$. While backpropagation through $M_{\psi}$ can be performed based on common first-order differentiation methods, a potential attacker must allow gradient flow through the modulation optimization steps performed in the forward pass of $R$. This requires differentiation through those optimization steps, namely via second-order differentiation and optimization. As we further show in Section 4.4, this may substantially increase the amount of computational resource required for the attacker.
The full PGD attack pipeline is portrayed in figure 2 - in each PGD iteration, the current perturbation is added the original signal. Then, the perturbation is kept frozen and the perturbed signal is being optimized for the corresponding INR $\theta^{*}$, that is being fed into the parameter-space classifier $M$.
3.3.2 TMO
Given the potentially immense computational resources required for full PGD attack optimization, we explore several attacks bypassing the need for full second-order differentiation in perturbation optimization – In Truncated Modulation Optimization (TMO) we apply a principal similar to the Truncated Backpropagation Through Time algorithm, widely used in RNN optimization. Since computational resources required for full PGD linearly grow with the number of modulation optimization steps $n$, in this attack we fix a certain parameter $\tau$ controlling the number of modulation optimization steps through which gradients would flow in attack optimization. Namely, rather than performing the forward pass of $R$ consisting of $n$ modulation optimization iterations, only $\tau$ optimization steps are performed - all of which will be backpropagated through.
Since the full number of modulation optimization steps $n$ is inherent within the attacked model, an adversarial perturbation optimized for a duration of $\tau$ modulation steps might fail to fool the classifier at inference time (even if the attacked succeeded after $\tau$ steps). To alleviate this gap, when optimizing an adversarial perturbation $\delta$ for a given data sample $x$, we first run the full $n$ optimization steps of $R(x)$ to find the ”clean” modulation vector. In every PGD iteration to follow, the $\tau$ optimization steps are initialized from the previously optimized ”clean” modulation.
3.3.3 BOTTOM
TMO (Section 3.3.2), while computationally light, suffers from the limitation of performing less modulation optimization step per every PGD iteration. Model behavior observed by the attacking algorithm therefore deviates from the full forward process of the attacked model. In Backpropagation Over Truncation Through Optimization of Modulation Optimization (BOTTOM) we address this, and propose an attack that performs the full $n$ modulation optimization steps per every sample $x$, while also alleviating the computational prices described in section 3.3.1. For every PGD iteration, the $n$ steps performed in the forward pass of $R$ are divided into $\left\lfloor\frac{n}{\tau}\right\rfloor$ optimization segments. In each segment, second order differentiation is performed through $\tau$ modulation optimization steps.
The attacker is given the flexibility to balance between computational resource consumption (lower $\tau$ values) and fidelity of gradients to the full forward pass in $R$ (higher $\tau$ values).
3.3.4 ICOP
Proposed attacks in previous sections focus on perturbations applied in signal domain. In this section we further propose an attack applied in INR domain. The main challenge here is efficiently constraining the adversarial perturbation - while we perturb the neural representation of the signal, we still must constrain the response of every applied perturbation in signal domain. Namely, unlike the formulation in equation 2, here we must impose $F(\delta)\leq B$. The need to query the Functa model $F$ to assess this response raises a difficulty to apply any trivial, immediate projection of a given perturbation $\delta$ onto the feasible set. We therefore propose Imposition of Constraints via Orthogonal Projection (ICOP) - this method relies on two approaches to impose soft constraints over the perturbations - First, after every PGD step we compute a projection loss, measuring constraint violation in $\mathcal{L}_{1}$: $\mathcal{L}_{proj.}=\|\mathcal{L}_{q}(F(\delta))-B\|_{1}$. We then iteratively take small gradient steps $\eta$: $\delta_{proj.}=\delta-\eta\frac{\partial\mathcal{L}_{proj.}}{\partial\delta}$, until constraint satisfaction.
Secondly, to discourage perturbation steps from severely breaking signal-domain constraints, we also found it useful to project the maximized classification loss to be orthogonal to $\frac{\partial\|F(\theta+\delta)-F(\theta)\|_{1}}{\partial\delta}$ (namely, the perturbation-induced deviation in signal-domain) in every perturbation gradient step. We enforce soft-constraints in optimization, and then impose hard constraints over the optimized perturbation to ensure feasibility. This is done by computing $R$ over $F(\delta)$ clamped onto the feasible set.
3.3.5 Implicit Differentiation
Implicit differentiation is a technique used to differentiate equations where the dependent and independent variables are intertwined, allowing for gradient computation in optimization problems where the input and output cannot be explicitly separated. In our adversarial setting (Equation 2), while $x$ and $M_{\psi}(R(x+\delta))$ are separable, any explicit differentiation method (e.g. computational graph-based gradient computation, as is the case in all previously proposed methods) would be heavily hindered by the nature of $R$. The successive optimization steps required for querying $R$ may vastly inflate the computational graph maintained for the adversarial objective, leading to increased computational costs and potential numerical instability commonly associated with large graphs.
To this end, we also investigate implicit differentiation within our setting. Let $d(x,\delta)$ denote the classification error (as featured in Equation 2), and $\mathscr{L}(\theta^{*},x)$ represent the INR signal approximation error $\|F(\theta^{*})-x\|_{\mathcal{X}}$.
Since $\theta^{*}=R(x)$, we elect the corresponding optimality conditions to represent local minima in $\theta^{*}$’s representation error of $x$:
$\frac{\partial\mathscr{L}(\theta^{*},x)}{\partial\theta^{*}}=0;\frac{\partial\mathscr{L}(\theta^{*}+d\theta,x+dx)}{\partial\theta^{*}}=0$
where $dx,d\theta$ represent infinitesimal steps in $x,\theta$ respectively. Our implicit differentiation is thus guided by the following approximation of gradients for our entire optimization pipeline:
$\frac{\partial d}{\partial\delta}=-\frac{\partial d}{\partial\theta^{*}}\cdot\frac{\partial^{2}\mathscr{L}}{\partial\theta^{*}\partial(x+\delta)}\cdot(\frac{\mathscr{L}}{\partial^{2}\theta^{*}})^{-1}\cdot\frac{\partial(x+\delta)}{\partial\delta}$ (3)
Note Equation 3 circumvents the evaluation of $\frac{\partial\theta^{*}}{\partial x}$ via the chain-rule , and therefore explicit backpropagation through the optimization steps applied when solving Equation 1 is not needed. Therefore, unlike previous methods proposed in this paper (Sections 3.3.1, 3.3.2, 3.3.3 and 3.3.4), the size of the computational graph remains constant with growth in the number of modulation optimization steps. In spite of its merits, we stress that implicit differentiation does not nullify the need of explicit methods – the efficacy of implicit differentiation is heavily reliant on the extent of satisfaction of the underlying optimality conditions. Specifically, in our setting, implicit differentiation is only reliable when the set of parameters $\theta^{*}$ given by $R(x)$ satisfies sufficient signal-domain fidelity with $x$.
3.4 BVA - Voxel-Grid Space Adversarial Attack
To study adversarial attacks over 3D structures, we must choose how to represent them, so that we can establish appropriate attack perturbation and constraints. Out of the 3 most common conventions for 3D data representations – voxel-grids, mesh-grids and point-clouds – in this work we specifically study adversarial perturbations over voxel-grid data. The reason for this choice is that our experiments show that due to the highly irregular structure of mesh-grids and point-clouds, fitting the dataset-shared set of INR weights ($\theta_{w}$ in notation from section 3.1) yields poor results, preventing achieving implicit representations of adequate signal fidelity.
While adversarial attacks over 3D have been studied in the past for mesh-grids ((Xu et al., 2022)) and point clouds (Liu et al., 2019a), to the best of our knowledge no similar work had so far been done in the context of voxel-grid representations. Since we wish to perform adversarial attacks over INR classifiers of voxel grids, and compare robustness to a corresponding signal-domain classifier, we develop Binary Voxel Attack (BVA) - a novel adversarial attack applied over voxel-grid data.
Voxel-grid representations are traditionally binary spatial occupancy maps. Therefore, an adversarial attack performing addition of a real-valued adversarial perturbation bounded in $\mathcal{L}_{\infty}$ is not applicable. Since our attacked signal is a binary map, we opt to establish an adversarial attack in the from of bit flipping (namely, bit attack). Given a voxel-grid input signal $x\in\{0,1\}^{H\times W\times D}$ (where $H,W,D$ are grid spatial dimensions), we formulate our attack as:
$\max_{\delta}\quad$ $\|M_{\psi}(x\oplus\delta)\|\text{s.t.}\|p\|_{0}=B$ (4)
where $\delta\in\{0,1\}^{H\times W\times D}$ is the optimized binary adversarial perturbation, and $\oplus$ denotes the bit-wise xor operation (performing bit switch over $x$ entries where $\delta$ has logical value of $1$). $\delta$ is chosen to be bounded in $\mathcal{L}_{0}$ norm, namely bounding the number of bits flipped in $x$.
This formulation calls for the incorporation of binary optimization techniques, as trivial binarization of $\delta$ (e.g. by thresholding) is non-differentiable. To optimize $\delta$ as a binary mask we follow the binary optimization technique presented in (Iliadis et al., 2020). This methods maintain two congruent masks - a binary mask $\Phi$ and a real mask $\Phi_{c}$. $\Phi$ is used for forward and backpropagation, where each gradient step applies the gradients calculated for $\Phi$ over $\Phi_{c}$. The latter is then binarized (without gradient tracking) to produce $\Phi$ for the next iteration.
To demonstrate inherent robustness of parameter-space classifiers over voxel-grid data, in section 4.2.1 we first exemplify the efficacy of signal-domain attacks using BVA. Then, in section 4.2.2 we employ BVA in parameter-space attacks and demonstrate increased adversarial robustness. We emphasize that equation 4 is a specific instance of equation 2, and therefore all parameter-space attacks previously proposed in section 3 can be performed on voxel-grids using BVA.
4 Results
In the following sections we demonstrate the inherent robustness of Parameter-Space classifiers over 3 different datasets - MNIST ((Deng, 2012)), Fashion-MNIST ((Xiao et al., 2017)) and ModelNet10 ((Wu et al., 2015)).
MNIST and Fashion-MNIST are chosen to represent the prevalent and heavily-studied modality of 2D images. While current adversarial robustness and computer-vision literature tends to focus on more challenging and diverse datasets (e.g. CIFAR100, ImageNet), applicability of these datasets in the parameter-space classification setting is currently limited. This is because state-of-the-art parameter-space classification algorithms proposed thus far have yet to achieve classification satisfying classification accuracies on such complex datasets ((Navon et al., 2023a, b; Chen & Wang, 2022)). ModelNet is used to demonstrate our findings over one of the modalities where usage of INRs is naturally useful, due to commonly large dimensionality. Following the approach from (Dupont et al., 2022), for each dataset $\mathcal{X}$ we first optimize a set of modulation vectors $\{v_{x}\forall x\in\mathcal{X}\}$. Then, we train a single MLP classifier over this modulation dataset. Finally, every pre-trained classifier is attacked with each of the five attacks presented in Section 3.3. We emphasize no robust are adversarial training is being done in any stage of this experimentation scheme. Attack formulations and results for 2D data are presented in section 4.1, and for 3D data in section 4.2. In section 4.4 we further demonstrate the added computational costs induced by performing adversarial attacks over deep weight-space classifiers, posing added practical challenges for potential attackers. Full training details are further specified in Section A.1.
4.1 Robustness For 2D Data
Clean and adversarial INR classification accuracy over 2D data classifiers is shown in table 1. In the notations of equation 2, $q=\infty$ is chosen as $B=\frac{\varepsilon}{256}$ for every table entry. Clean column represents classifier accuracy without introducing any adversarial perturbations.
Results indicate 3 major trends - First, both MNIST and Fashion-MNIST are adversarially robust. For MNIST, while the TMO and BOTTOM attacks are the only two that managed to incur any above-marginal decrease in classification accuracy, it is still far below the accuracy decrease of around $60\%$ that is traditionally observed in similarly-bounded PGD attacks over signal-domain classifiers without any robust training ((Wong et al., 2020)). While higher-bounded attacks over the Fashion-MNIST dataset have been more successful, we claim their success is substantially lower than the effect of a similarly constrained attack over a signal-domain classifier. To show this, we train four baseline signal-domain classifiers and attack with PGD using similar $\mathcal{L}_{\infty}$ norm constraints and PGD iteration number. We use Vision-Transformers (ViTs) (Dosovitskiy, 2020) due to their outstanding record in natural image classification (Han et al., 2022), and two convolutional models for their ubiquity and foundational role in computer-vision. To experiment with models of different sizes we train a ViT-Large model and a Resnet18, as well as a ViT-Base model and a compact,2-layer CNN. Full architecture details are included in Section A.3.2.
We compare the robust accuracy of these four Signal-space classifiers under standard 100-iteration PGD attack, with that of our parameter-space classifier subjected to all attacks from Sections 3.3.1, 3.3.2, 3.3.3, 3.3.4 and 3.3.5. Results, shown in figure 3, show parameter-space classifiers show significantly higher robustness, especially for lower attack bounds.
The second indicated trend is that we observe that the full PGD attack demonstrates lower adversarial prowess compared to BOTTOM and TMO attacks in all experiments. We attribute this to the fact that in full-PGD the numerous steps of second-order optimization required for standard PGD in our case encourage vanishing gradients phenomena, namely attenuating gradient information backpropagated through the modulation optimization process $R$. These difficulties are added to the computational encumbrance of running full PGD (see section 4.4), and further stress the need in initial methods developed here (sections 3.3.2&3.3.3), as well as further related research. While computationally lightest, implicit differentiation proves less potent than most explicit methods. We attribute this to the strong dependence of this method in fulfillment of specified optimality conditions.
Lastly, we conclude that the INR-domain attack (ICOP) failed to impose any actual effect on classification accuracy. We attribute this result to the heuristic nature of the approaches used to constrain the INR-domain perturbation’s response in the signal domain (section 3.3.4). We still include it within our suite of proposed attacks, as it proved more efficient when initiated over 3D data (section 4.2).
Figure 3: Robust accuracy under $\mathcal{L}_{\infty}$ norm constraints - for parameter-space (solid curves) and signal-space (dotted curves) classifier attacks.
Table 1: MNIST & Fashion-MNIST Clean/Robust Classification Accuracy - for any $\varepsilon$ value, $\mathcal{L}_{\infty}$ attack bound (equation 2) is $B=\frac{\varepsilon}{256}$.
$\varepsilon$ MNIST Fashion-MNIST
Clean Full TMO BOTTOM ICOP Implicit Clean Full TMO BOTTOM ICOP Implicit
4 0.976 0.949 0.941 0.939 0.965 0.957 0.891 0.791 0.77 0.79 0.891 0.826
8 0.976 0.915 0.91 0.906 0.965 0.946 0.891 0.67 0.64 0.658 0.8904 0.792
16 0.976 0.866 0.825 0.821 0.961 0.926 0.891 0.54 0.44 0.438 0.89 0.745
32 0.976 0.803 0.7035 0.698 0.957 0.919 0.891 0.48 0.37 0.341 0.848 0.69
4.2 Robustness For 3D Data
In section 4.1 we survey adversarial robustness of classifiers INR for 2D data. The importance of this baseline is comparability to the widely studied adversarial robustness of 2D classifiers. Nonetheless, practical advantages of using INRs for 2D data are usually limited, since common grid representations suffice for most common tasks. Data modalities of larger dimensions, such as 3D data, however, more commonly gain from using INRs for achieving downstream tasks, including classification ((Jiang et al., 2020; Molaei et al., 2023; Niemeyer et al., 2020)). We use the ModelNet 10 dataset ((Wu et al., 2015)) consisting of 4900 3D CAD models given in Mesh-grid representation.
4.2.1 Signal Space Attack
To demonstrate inherent adversarial robustness over 3D data under the BVA attack (section 3.4), we must first demonstrate its prowess in attenuating signal-domain classification accuracy. To do so we train a voxel-grid classification model based on Point-Voxel CNN layers proposed by (Liu et al., 2019b). We train the classifier over ModelNet10, converting every mesh-grid sample to a voxel-grid of dimension $15\times 15\times 15$ using Trimesh ((Dawson-Haggerty et al., )). Then, we attack this trained model with BVA for a wide range of $\mathcal{L}_{0}$ bounds and assess attacked classification accuracy. Results are displayed in figure 4, showing BVA is efficient over the tested signal-space classifier, inducing exponential decay of classification accuracy with growth in number of flipped voxel bits.
Figure 4: BVA robust accuracy for signal-space classifier and weight-space classifier - under $\mathcal{L}_{0}$ norm constraints.
4.2.2 Parameter Space Attack
Figure 4 compares the robust accuracy of the signal domain (black) and parameter-space classifiers, where the parameter-space classifier is attacked with TMO, BOTTOM, ICOP, and implicit attacks. We note that we do not experiment with the full PGD attack in parameter-space for the 3D INR classifier, since to optimize INR for the ModelNet10 dataset with sufficient input signal fidelity, we had to perform 500 modulation optimization iterations per sample. While computation times for doing so are not very high (several seconds), backpropagation through these 500 optimization steps calls for infeasible memory and compute time requirements, and also lead to severe vanishing/exploding gradient issues. We further elaborate on these considerations in Section 4.4.
While the parameter-space classifier exhibits lower clean accuracy (73% in parameter-space against 82% in signal space), results demonstrate conclusive increased adversarial robustness for the signal-space classifier. We can specify two major trends - the first is that the absolute decrease in classification accuracy is much lower for the parameter-space classifier. The second is that for the attacked weight-space classifier, the decrease in accuracy attenuation across bound increase saturates significantly faster. The data from Figure 4 is presented numerically in Table 2.
4.3 Qualitative Results
In this section we provide qualitative evidence for the inherent robustness of parameter-space classifiers. Adversarial attacks are known to effectively exploit the amplifications of small input-space perturbations (Goodfellow, 2014; Haleta et al., 2021). As previously discussed, we argue that the enhanced robustness of parameter-space classifiers arises from the attacker’s inability to directly perturb the classifier’s input. Instead, the attacker must optimize a constrained signal-domain adversarial perturbation for which the response received as input to the classifier (after modulation optimization Equation 1) is adversarially potent. This is deemed to be substantially more challenging for the attacker, as modulation optimization tends to attenuate adversarial prowess fed as input to the classifier.
To illustrate this phenomenon, we report the mean amplification, measured as the difference in activation values between perturbed and clean input data, for each layer in our pipeline (Figure 2). These measurements, shown in Figure 5, were calculated using the MNIST dataset, with every sample propagated through the pipeline after being perturbed by each attack in our suite with a constraint of $\varepsilon=\frac{32}{256}$. To further validate the relationship between attack success and amplification, we differentiate between robust and non-robust samples. Successful attacks demonstrate an average amplification approximately three times greater than that of failed attacks. In both cases, amplification sharply plummets upon entering the first classifier layer (layer 10) and remains consistently low in subsequent layers.
These findings indicate that adversarial amplification within the INR layers has a negligible impact on the downstream classification model. This supports our claim that modulation optimization effectively obfuscates the adversarial potency of perturbations applied in the signal domain. Importantly, this effect is not observed for signal-domain classifiers, as we show in Figure 6. Results for other attack constraints and datasets are reported in Appendix B.
Figure 5: Adversarial amplification across pipeline layers - showing modulation optimization’s attenuation of adversarial amplification.
In Figure 1 we demonstrate the affect of adversarial perturbations applied in signal-domain over the latent representations attained from perturbed data, and compare between parameter-space classifiers and traditional signal-space classifiers. To this end, as a signal-domain baseline, we first train a convolutional autoencoder over the MNIST dataset. For a fair comparison, this autoencoder is trained disjointly from the downstream classification task. We then train a classifier over the learned latent space, and finally perform 100 iterations of signal-domain adversarial attack over this classifier. The bottom half of Figure 1 illustrates t-SNE projection of the trained autoencoder’s latent representations before (left) and after (right) the adversarial attack. The top half of the plot corresponds to parameter-space classifiers, showing t-SNE projections of INR modulations optimized for clean (right) data and for data attacked with our full-PGD attack (Section 3.3.1. Prior to the attacks, the latent space of both signal and parameter-space domains sis well-structured, with clear separation between the different classes. However, after the attack, the model operating strictly in signal-domain (i.e. the autoencoder baseline) experiences severe disruption of this structure – classes of adversarial samples within the classifier’s latent input space are no longer distinct and are instead intermixed. This indicates a significant deterioration in the class separability, highlighting the impact of the adversarial perturbations on the latent space. Adversarial attacks optimized for the parameter-space classifier, on the other hand, seem inconsequential to the input-space structure of the parameter-space classifier, including class separability. These conclusions are well-correlated with those from Figure 5.
4.4 Computational Resource Analysis
In sections 4.1,4.2 we provide evidence for the adversarial robustness of parameter-space classifiers based on classification accuracies under the presence of adversarial perturbations. In this section, we further support these claims from a different, practical perspective. We show that orthogonally to the efficacy of adversarial attacks over parameter-space classifiers, attacking these classifiers also poses significant added computational difficulties for the potential attacker. These difficulties stem from the optimization process inherent within the classification pipeline’s inference process. To illustrate this, we measure the computation times required for a single modulation optimization step during a single PGD iteration on an NVIDIA RTX-2080 GPU. Our results show that for the MNIST dataset with 512-dimensional modulation vectors, the compute time per step is approximately 0.015 seconds, while for the ModelNet10 dataset with 2048-dimensional modulation vectors, it is 0.02 seconds. Even for a small number of PGD steps, increasing the number of modulation optimization steps – an aspect controlled by the model, not the attacker – increases attack optimization times within several orders of magnitude. Importantly, this increase does not significantly impact the computation time required for ”clean” (unperturbed) inference. For instance, with 100 modulation optimization steps set by the model to solve Equation 1, ”clean” inference would require 1.5 seconds per sample, while perturbation optimization of 100 PGD iterations would require the attacker to invest 150 seconds per-sample.
On top of computation time increase, attacking parameter-space classifiers via explicit differentiation also poses a challenge for the attacker in terms of memory consumption. Tested on the same hardware, we find that GPU memory consumption for the full PGD attack over the ModelNet10 classifier grows linearly with number of modulation optimization steps, with a slope of $0.1705[\frac{GB}{Mod.Steps}]$. The corresponding coefficient for the smaller MNIST data is $0.049[\frac{GB}{Mod.Steps}]$. Namely, in order to perform a full PGD (Section 3.3.1) attack over the ModelNet10 data, every additional modulation optimization step included in ”clean” model training induces roughly 170 additional megabytes for the attacker’s backpropagation. Backpropagating through the 500 modulation optimization steps used in creation of the ModelNet10 INR dataset in our experiments would require 85 Gigabytes of GPU memory, greatly increasing optimization costs. The increase in modulation optimization steps holds no more than marginal prices for the ”innocent” (non-attacker) user in INR dataset optimization, as this process does not require 2nd-order differentiation. Implicit differentiation (Section 3.3.5) offers the attacker constant-memory attacks (in our experiments - 526 MB for MNIST, 986 MB for ModelNet10), however at the cost of higher sensitivity to optimality condition satisfaction (as shown in Section 4.1).
5 Discussion
5.1 Conclusion
In this work, we’ve explored the adversarial robustness of parameter-space classifiers. We’ve presented the challenges involved in translating common PGD-based attacks onto models performing classification directly in deep parameter-space, and addressed these challenges by developing a novel set of five different adversarial attacks designed for parameter-space classification, applying perturbations both in the signal domain and in parameter space. We then tested these attack over three different datasets from two popular modalities, and shown that parameter-space classifiers exhibit an increased inherent adversarial robustness, compared to their counterparts classifying in signal-domain. We’ve further presented the notion of ”robustness” stemming from the potentially large added computational resources required for certain cases of parameter-space classifier attacks.
5.2 Limitations and Future Work
While we consider our work an important initial study of the currently under-explored topic of adversarial robustness in parameter-space classification, several notable limitations remain. First, in this work we consider a relatively limited set of simple datasets from both 2D and 3D modalities. This choice mainly stems from the fact that current state-of-the-art parameter space classifiers have yet to achieve adequate ”clean” (non-attacked) classification results over more challenging datasets ((Navon et al., 2023a, b)). Since the focus of our work is adversarial robustness of parameter-space classifiers, experimenting with our proposed attacks over classifiers with low ”clean” accuracy would make assessment of our work less reliable. We hope future exploration of methods for better non-adversarial parameter-space classification would expand our set of potential datasets for robustness exploration. Second, we note that the adversarial robustness potentially attained when using parameter-space classifiers often comes with the cost of reduced classification accuracy, as shown in section 4. Finally, our work focuses on inherent adversarial robustness, and does not discuss active defenses. We deem our findings about inherent robustness an encouraging starting point for future research in development of adversarial defenses and robust modeling.
Appendix A Training Regime, Hyperparameters & Technical Details
A.1 Functaset Generation
A.1.1 MNIST & Fashion-MNIST
All experiments described in this section have been conducted on a single NVIDIA RTX-2080 GPU. Following (Dupont et al., 2022), for each of the three datasets considered in this paper, we first train a SIREN (Sitzmann et al., 2020) jointly over the entire training set. For MNIST and Fashion-MNIST we optimize modulation vectors of dimension $512$ for 6 epochs of 3 modulation optimization steps per-sample, with a learning-rate of $0.000005$ and a batch size of $128$. We use 10 SIREN layers with hidden dimension of $256$. We then proceed to creating the Functaset , which is the dataset of modulation vectors conditioning the SIREN for per-sample representation. For each sample we optimize for 10 modulation optimization steps, with a learning-rate of 0.01.
A.1.2 ModelNet10
For ModelNet10 we similarly follow (Dupont et al., 2022), however we first convert the Mesh-grid data onto a voxelgrid, as previously mentioned (Section 3.4). Each voxelgrid is resampled to dimension $15\times 15\times 15$. Due to the increased dimensionality, we train $2048$-dimensional Functa modulations with a SIREN similar to the one used for the MNIST dataset. After training the shared SIREN weights, each modulation vector is further optimized for $500$ modulation optimization steps.
A.2 Adversarial Attacks
After the Functaset is generated, we proceed to executing our develop suite of adversarial attacks Section 3.3. For MNIST and Fashion-MNIST, all attacks are executed over 10 modulation optimization steps per adversarial perturbation, similarly to optimization in the clean case. BOTTOM and TMO optimization segments are chosen at 5 steps per-sample. For implicit differentiation, we’ve found it most useful to optimize each perturbed image using L-BFGS, rather than gradient methods (which have been used in other attacks). Each L-BFGS optimization is carried for at most 20 iterations. For ModelNet10, ICOP,TMO and BOTTOM attacks are performed for a total of 500 modulation optimization steps per perturbed voxelgrid, with attack segments again chosen at 5 steps. Implicit differentiation attack is also performed with at most 20 iterations of L-BFGS. All attacks are carried-out for 100 PGD iterations, with early-stop of 3 iterations of constant loss for the implicit method. We use the Adam optimizer in all experiments.
A.3 Baseline Models
A.3.1 Classifier Baselines
In this section we specify implementation and training details related to the signal-space classifier baseline CNN and ViT architectures used in Section 4. In section 4.1 we train 4 signal-domain classifiers – a small CNN, a ResNet18, ViT-Large/16 and Vit-Base/16. For the ViTs we leverage pre-trained models from (Wightman, 2019), and finetune for our classification task using a a linear probe prior to a applying Softmax. For the Resnet we train from scratch, and project encoded data with a linear probe followed by a Softmax. For the small-CNN we use two convolutional layers with ReLU activations, and a single MaxPooling operation between them. For all baselines we train over the Fashion-MNIST training set for 5 epochs with learning-rate of 0.0001.
A.3.2 Autoencoder Baselines
In Section 4.3, we present a qualitative analysis of adversarial perturbation amplification across pipeline layers, along with t-SNE projections of modulation spaces before and after adversarial perturbation. To connect these observations to the inherent adversarial robustness of parameter-space classifiers, we include a comparison with a signal-space classifier. In order to propose a signal-space classifier with an appropriate counterpart to the explored parameter-space model modulation space, we choose to train classifiers operating over latent spaces trained by autoencoders. We model each autoencoder as a 3-layer CNN with ReLU activations and train with a self-supervised reconstruction objective for 10 epochs. After the autoencoder is trained, we freeze it and train a linear classifier with 2 hidden layers and a single ReLU activation on top of it. The frozen encoder layers acts as the INR counterpart in Figure 5, and the MLP head is parallelized to the parameter-space classifier. We use a latent dimension of $128$ for all encoders.
Appendix B Additional Results
B.1 Robustness For 2D Data
B.1.1 MNIST
Figure 6 illustrates the mean amplification difference measured through model activation layers over MNIST data, similarly to Figure 5, however for the signal-domain classifier trained according to section A.3.2. This model uses a 3-layer convolutional encoder, followed by a classifier with a single activation. Clean classifier accuracy is $98.8\%$, robust accuracy is $11.28\%$. Differently from similar results displayed for parameter-space classifiers in Figure 5, upon transition from encoder to classifier layers (3rd layer, marked by the black vertical line), the adversarial amplification is not attenuated, but rather drastically inflated.
Figure 6: Signal-Domain Amplification - through encoder and downstream classifier layers
Figure 7 showcases adversarial amplification (similarly to Figure 5) across different attack bound for the full-PGD attack. Adversarial obfuscation drawn from Figure 5 arises here as well.
Figure 7: Signal-Domain Amplification - through encoder and downstream classifier layers, across different constraints $\frac{\varepsilon}{256}$ for full PGD attack
B.1.2 Fashion-MNIST
Figure 8 illustrates adversarial amplification across pipeline layers, similarly to Figure 5. Trends of correlation between amplification and adversarial success, as well as evident obfuscation of adversarial affect between INR and classifier segments, are consistent with findings from Section 4.3.
Figure 8: Adversarial amplification across pipeline layers - showing modulation optimization’s attenuation of adversarial amplification.
Similarly to Figure 6, we test amplification patterns of a signal-domain classifier (Section A.3.2) for the Fashion-MNIST dataset as well. We again observe consistency across both datasets - the signal domain classifiers proves to be much more sensitive to adversarial examples propagated through the encoder-classifier pipeline.
Figure 9: Signal-Domain Amplification - through encoder and downstream classifier layers
Figure 10 exhibits the affect of adversarial attacks in signal-domain over the underlying modulation vectors (for parameter-space classifiers) and latent vectors (for signal-space classifiers) received as classifier input. Similarly to Figure 1, for both parameter-space and signal-space model, a well-formed latent space is observed before the incorporation of adversarial perturbation. After attacking with PGD, the corresponding modulation vectors fed into the parameter-space classifier remain almost unchanged, whilst the latent space observed by the signal-domain classifier is heavily deformed.
Figure 10: t-SNE projection of modulation vectors fitted for clean and adversarialy-perturbed data - for parameter-space and signal-space classifiers.
B.2 Robustness For 3D Data
In this section we report additional results for the ModelNet10 dataset.
In Table 2 we report numerical results for the data from Figure 4 The first two columns represent the same attack constraints in absolute bit numbers ($\varepsilon$) and bit numbers relative to the entire volume (matching horizontal axis in figure 4).
Table 2: ModelNet10 Clean/Robust Classification Accuracy - for any given $\varepsilon$ value, attack bound from equation 2 is $B=\varepsilon$ in $\mathcal{L}_{0}$ (i.e. number of flipped signal bits).
$\varepsilon$ Flip Ratio Clean TMO BOTTOM ICOP Implicit
10 0.003 0.734 0.669 0.663 0.66 0.722
20 0.006 0.734 0.6475 0.634 0.6475 0.708
40 0.012 0.734 0.647 0.623 0.641 0.696
80 0.024 0.734 0.643 0.616 0.581 0.67
100 0.03 0.734 0.6412 0.609 0.557 0.657
200 0.06 0.734 0.594 0.576 0.41 0.613

### v2

@@
-Adversarial Robustness in Parameter-Space Classifiers
+Adversarial Robustness in Parameter-Space Classifiersthanks: Code available at https://github.com/tamirshor7/Parameter-Space-Attack-Suite
 Abstract
@@
 To support our claims, we develop a novel suite of adversarial attacks targeting parameter-space classifiers, and furthermore analyze practical considerations of attacking parameter-space classifiers.
-Code for reproducing all experiments and implementation of all proposed methods will be released upon publication.
 1 Introduction
 Implicit Neural Representations (INRs) are representations of an arbitrary signal as a neural network predicting signal values under some conditioning.
-These representations have recently been raising increased interest due to their ability to efficiently and compactly encode high-dimensional data while maintaining high input fidelity, and have further shown to be capable of learning additional intrinsic knowledge latent within the original signal-domain representation (Sitzmann et al., 2020; Wang et al., 2021).
-Figure 1: t-SNE projection of modulation vectors fitted for clean and adversarialy-perturbed data - for parameter-space and signal-space classifiers.
-While INRs are inherently designed to fit one specific signal (rather than generalizing across an entire dataset), several recent works have proposed ways to use these representations to achieve a variety of downstream tasks by using some neural meta-network operating directly over the parameter-space of the INR of each signal within the dataset ((Chen & Wang, 2022; Dupont et al., 2022; Lee et al., 2021; Navon et al., 2023b)).
+These representations have recently been raising increased interest due to their ability to efficiently and compactly encode high-dimensional data while maintaining high input fidelity, and have further shown to be capable of learning additional intrinsic knowledge latent within the original signal-domain representation Sitzmann et al.
+(2020); Wang et al.
+(2021).
+While INRs are inherently designed to fit one specific signal (rather than generalizing across an entire dataset), several recent works have proposed ways to use these representations to achieve a variety of downstream tasks by using some neural meta-network operating directly over the parameter-space of the INR of each signal within the dataset (Chen & Wang (2022); Dupont et al.
+(2022); Lee et al.
+(2021); Navon et al.
+(2023b)).
 In these settings, each data sample is received in the native signal space, converted to its compressed INR representation by fitting a neural model, and finally the meta-network performs the designated downstream task over the implicit representation encoded by the model’s parameters.
@@
 We then proceed to show that discussed parameter-space classification models are inherently robust to adversarial attacks in comparison to the signal-domain classification problem counterpart, as exemplified using our established attack suites.
-Our hypothesis of inherent robustness can be motivated by the following: Firstly, adversarial attacks commonly utilize the highly-nonlinear nature of modern deep classifiers to create perturbations with specific patterns that, while bounded, will be heavily amplified by the classifier to create significant data manipulations in deeper layers ((Ingle & Kulkarni, 2021; Romano et al., 2020)).
+Our hypothesis of inherent robustness can be motivated by the following: Firstly, adversarial attacks commonly utilize the highly-nonlinear nature of modern deep classifiers to create perturbations with specific patterns that, while bounded, will be heavily amplified by the classifier to create significant data manipulations in deeper layers (Ingle & Kulkarni (2021); Romano et al.
+(2020)).
 In this setting of parameter-space classification, INR optimization is performed per-sample, and is embedded onto the classifier’s inference process.
 This constrains the attacker to optimize perturbations in the signal-domain rather than directly manipulating the classifier’s input.
-This optimization is disjoint from the downstream objective, and is dedicated to expressing global data features as a whole, rather than entangle specific entries with specific local signal features (Sitzmann et al., 2020).
+This optimization is disjoint from the downstream objective, and is dedicated to expressing global data features as a whole, rather than entangle specific entries with specific local signal features Sitzmann et al.
+(2020).
 Therefore, this optimization process tends to obfuscate adversarial patterns introduced in signal domain.
-Secondly, as we later show in Section 4.4, performing adversarial attacks over parameter-space classifiers poses significant computational difficulties for a white-box attacker, as attacking the INR given a perturbed signal requires backpropagation through an internal optimization loop.
-In summary, this paper makes the following contributions:
-  1. 1.
-    We supply evidence of inherent robustness of parameter-space classifiers to adversarial attacks.
-  2. 2.
-    We propose several novel types of adversarial white-box attacks for parameter-space classifiers.
-  3. 3.
-    To demonstrate our claims on 3D data, we also formulate a novel adversarial attack for voxel-grid data representations, and demonstrate its performance in both parameter-space and signal-space classification settings.
+Secondly, performing adversarial attacks over parameter-space classifiers poses significant computational difficulties for a white-box attacker, as attacking the INR given a perturbed signal requires backpropagation through an internal optimization loop.
+In summary, this paper makes the following contributions – First, we supply evidence of inherent robustness of parameter-space classifiers to adversarial attacks.
+Second, we propose several novel types of adversarial white-box attacks for parameter-space classifiers.
+Lastly, to demonstrate our claims on 3D data, we also formulate a novel adversarial attack for voxel-grid data representations.
 2 Related Work
 The analysis of trained neural network parameters had been raising increasing interest in recent years, demonstrating applicability in a wide, and at times surprising, range of machine learning tasks.
-(Eilertsen et al., 2020) was one of the first works considering this idea, performing trained network parameter analysis to identify the optimization procedure hyperparameters.
-(Unterthiner et al., 2020) showed that a trained CNN classifier model parameters can be used to directly attain a good assessment of classification accuracy, without running the full model’s forward pass.
-(Dupont et al., 2022) have been the first to demonstrate that neural network parameters, specifically of the SIREN (Sitzmann et al., 2020) models (one of the first INRs to be proposed), could be used as the direct and only input for the training of a meta-learner model for classification, generative, data-imputation and novel-view synthesis tasks.
-(Peebles et al., 2022) employ generative modeling of deep-weight space to improve parameter update steps in model training.
+Eilertsen et al.
+(2020) was one of the first works considering this idea, performing trained network parameter analysis to identify the optimization procedure hyperparameters.
+Unterthiner et al.
+(2020) showed that a trained CNN classifier model parameters can be used to directly attain a good assessment of classification accuracy, without running the full model’s forward pass.
+Dupont et al.
+(2022) have been the first to demonstrate that neural network parameters, specifically of the SIREN Sitzmann et al.
+(2020) models (one of the first INRs to be proposed), could be used as the direct and only input for the training of a meta-learner model for classification, generative, data-imputation and novel-view synthesis tasks.
+Peebles et al.
+(2022) employ generative modeling of deep-weight space to improve parameter update steps in model training.
 3 Method
@@
 We denote a parameter-space classifier as $M_{\psi}$, parameterized by $\psi$; the model receives the representation parameters $\theta=R(x)$ as the input and outputs the predicted class label $\hat{c}=M_{\psi}(\theta)$ as the output.
-Even for small neural representation models, the parameter space $\mathcal{Y}$ is often high-dimensional, increasing difficulties related to computational resources and overfitting when training the classification models on it (Navon et al., 2023a, b).
+Even for small neural representation models, the parameter space $\mathcal{Y}$ is often high-dimensional, increasing difficulties related to computational resources and overfitting when training the classification models on it (Navon et al., 2023a; b).
 Functas (Dupont et al., 2022) cope with this issue of increased parameter-space dimensionality by partitioning the parameter space into weights and biases, $\mathcal{Y}=\mathcal{Y}_{\mathrm{w}}\cup\mathcal{Y}_{\mathrm{b}}$, and sharing a single vector of weights $\theta_{\mathrm{w}}\in\mathcal{Y}_{\mathrm{w}}$ across a collection of represented signals, while allowing to fit individual modulation vectors $v(x)\in\mathbb{R}^{d}$ of a fixed dimension with a learned shared linear decoder $W$ such that each signal has an individual vector of biases, $\theta_{\mathrm{b}}(x)=Wv(x)$.
-This technique was successfully demonstrated with the SIREN (Sitzmann et al., 2020) models.
+This technique was successfully demonstrated with the SIREN (Sitzmann et al.
+(2020)) models.
 In what follows, we continue denoting the representation of $x$ as $\theta$, while taking into account that the actual INR $F$ may have additional parameters shared across multiple signals.
@@
 While backpropagating through this optimization process is possible (e.g.
-via automatic differentiation or implicit differentiation (Navon et al., 2020)), as we later show in Section 4, doing so dramatically increases computational resources required for perturbation optimization, thus posing additional practical challenges for a potential attacker, in addition to the inherent difficulty of optimization such an attack that stems from the inference pipeline itself, as motivated in section 1.
+via automatic differentiation or implicit differentiation Navon et al.
+(2020)), as we later show in Section 4, doing so dramatically increases computational resources required for perturbation optimization, thus posing additional practical challenges for a potential attacker, in addition to the inherent difficulty of optimization such an attack that stems from the inference pipeline itself, as motivated in section 1.
 In the following sections, we propose five different adversarial attacks, addressing these challenges that are absent in the traditional signal-domain adversarial attack settings.
@@
 This requires differentiation through those optimization steps, namely via second-order differentiation and optimization.
-As we further show in Section 4.4, this may substantially increase the amount of computational resource required for the attacker.
+As we further show in appendix C, this may substantially increase the amount of computational resource required for the attacker.
 The full PGD attack pipeline is portrayed in figure 1 - in each PGD iteration, the current perturbation is added the original signal.
@@
 The need to query the Functa model $F$ to assess this response raises a difficulty to apply any trivial, immediate projection of a given perturbation $\delta$ onto the feasible set.
-We therefore propose Imposition of Constraints via Orthogonal Projection (ICOP) - this method relies on two approaches to impose soft constraints over the perturbations - First, after every PGD step we compute a projection loss, measuring constraint violation in $\mathcal{L}_{1}$: $\mathcal{L}_{proj.}=\|\mathcal{L}_{q}(F(\delta))-B\|_{1}$.
+We therefore propose Imposition of Constraints via Orthogonal Projection (ICOP) - after every PGD step we compute a projection loss, measuring constraint violation in $\mathcal{L}_{1}$: $\mathcal{L}_{proj.}=\|\mathcal{L}_{q}(F(\delta))-B\|_{1}$.
 We then iteratively take small gradient steps $\eta$: $\delta_{proj.}=\delta-\eta\frac{\partial\mathcal{L}_{proj.}}{\partial\delta}$, until constraint satisfaction.
-Secondly, to discourage perturbation steps from severely breaking signal-domain constraints, we also found it useful to project the maximized classification loss to be orthogonal to $\frac{\partial\|F(\theta+\delta)-F(\theta)\|_{1}}{\partial\delta}$ (namely, the perturbation-induced deviation in signal-domain) in every perturbation gradient step.
-We enforce soft-constraints in optimization, and then impose hard constraints over the optimized perturbation to ensure feasibility.
-This is done by computing $R$ over $F(\delta)$ clamped onto the feasible set.
+To discourage perturbation steps from severely breaking signal-domain constraints, we also found it useful to project the maximized classification loss to be orthogonal to the perturbation-induced deviation in signal-domain $\frac{\partial\|F(\theta+\delta)-F(\theta)\|_{1}}{\partial\delta}$ in every perturbation gradient step.
 3.3.5 Implicit Differentiation
-Implicit differentiation is a technique used to differentiate equations where the dependent and independent variables are intertwined, allowing for gradient computation in optimization problems where the input and output cannot be explicitly separated.
-In our adversarial setting (Equation 2), while $x$ and $M_{\psi}(R(x+\delta))$ are separable, any explicit differentiation method (e.g.
+In our adversarial setting (eq.
+2), while $x$ and $M_{\psi}(R(x+\delta))$ are separable, any explicit differentiation method (e.g.
 computational graph-based gradient computation, as is the case in all previously proposed methods) would be heavily hindered by the nature of $R$.
@@
 To this end, we also investigate implicit differentiation within our setting.
-Let $d(x,\delta)$ denote the classification error (as featured in Equation 2), and $\mathscr{L}(\theta^{*},x)$ represent the INR signal approximation error $\|F(\theta^{*})-x\|_{\mathcal{X}}$.
+Let $d(x,\delta)$ denote the classification error (as featured in eq.
+2), and $\mathscr{L}(\theta^{*},x)$ represent the INR signal approximation error $\|F(\theta^{*})-x\|_{\mathcal{X}}$.
 Since $\theta^{*}=R(x)$, we elect the corresponding optimality conditions to represent local minima in $\theta^{*}$’s representation error of $x$:
@@
 $\frac{\partial d}{\partial\delta}=-\frac{\partial d}{\partial\theta^{*}}\cdot\frac{\partial^{2}\mathscr{L}}{\partial\theta^{*}\partial(x+\delta)}\cdot(\frac{\mathscr{L}}{\partial^{2}\theta^{*}})^{-1}\cdot\frac{\partial(x+\delta)}{\partial\delta}$ (3)
-Note Equation 3 circumvents the evaluation of $\frac{\partial\theta^{*}}{\partial x}$ via the chain-rule , and therefore explicit backpropagation through the optimization steps applied when solving Equation 1 is not needed.
-Therefore, unlike previous methods proposed in this paper (Sections 3.3.1, 3.3.2, 3.3.3 and 3.3.4), the size of the computational graph remains constant with growth in the number of modulation optimization steps.
+Note that unlike previous methods proposed in this paper (sections 3.3.1, 3.3.2, 3.3.3 and 3.3.4), eq.
+3 circumvents the evaluation of $\frac{\partial\theta^{*}}{\partial x}$ via the chain-rule , and therefore explicit backpropagation through the optimization steps applied when solving eq.
+1 is not needed.
 In spite of its merits, we stress that implicit differentiation does not nullify the need of explicit methods – the efficacy of implicit differentiation is heavily reliant on the extent of satisfaction of the underlying optimality conditions.
-Specifically, in our setting, implicit differentiation is only reliable when the set of parameters $\theta^{*}$ given by $R(x)$ satisfies sufficient signal-domain fidelity with $x$.
+Under our formulation – when the set of parameters $\theta^{*}$ given by $R(x)$ satisfies sufficient signal-domain fidelity with $x$.
 3.4 BVA - Voxel-Grid Space Adversarial Attack
-To study adversarial attacks over 3D structures, we must choose how to represent them, so that we can establish appropriate attack perturbation and constraints.
-Out of the 3 most common conventions for 3D data representations – voxel-grids, mesh-grids and point-clouds – in this work we specifically study adversarial perturbations over voxel-grid data.
-The reason for this choice is that our experiments show that due to the highly irregular structure of mesh-grids and point-clouds, fitting the dataset-shared set of INR weights ($\theta_{w}$ in notation from section 3.1) yields poor results, preventing achieving implicit representations of adequate signal fidelity.
-While adversarial attacks over 3D have been studied in the past for mesh-grids ((Xu et al., 2022)) and point clouds (Liu et al., 2019a), to the best of our knowledge no similar work had so far been done in the context of voxel-grid representations.
+Out of the 3 most common conventions for 3D data representations – voxel-grids, mesh-grids and point-clouds – we find that due to the highly irregular structure of mesh-grids and point-clouds, fitting the dataset-shared set of INR weights ($\theta_{w}$ in notation from section 3.1) yields poor results, preventing achieving implicit representations of adequate signal fidelity.
+Therefore, in this work we focus on voxel-grid representations of 3D data.
+While adversarial attacks over 3D have been studied in the past for mesh-grids (Xu et al.
+(2022)) and point clouds Liu et al.
+(2019a), to the best of our knowledge no similar work had so far been done in the context of voxel-grid representations.
 Since we wish to perform adversarial attacks over INR classifiers of voxel grids, and compare robustness to a corresponding signal-domain classifier, we develop Binary Voxel Attack (BVA) - a novel adversarial attack applied over voxel-grid data.
@@
 $\max_{\delta}\quad$ $\|M_{\psi}(x\oplus\delta)\|\text{s.t.}\|p\|_{0}=B$ (4)
-where $\delta\in\{0,1\}^{H\times W\times D}$ is the optimized binary adversarial perturbation, and $\oplus$ denotes the bit-wise xor operation (performing bit switch over $x$ entries where $\delta$ has logical value of $1$).
+where $\delta\in\{0,1\}^{H\times W\times D}$ is the optimized binary adversarial perturbation, and $\oplus$ denotes the bit-wise xor operation.
 $\delta$ is chosen to be bounded in $\mathcal{L}_{0}$ norm, namely bounding the number of bits flipped in $x$.
-This formulation calls for the incorporation of binary optimization techniques, as trivial binarization of $\delta$ (e.g.
-by thresholding) is non-differentiable.
-To optimize $\delta$ as a binary mask we follow the binary optimization technique presented in (Iliadis et al., 2020).
-This methods maintain two congruent masks - a binary mask $\Phi$ and a real mask $\Phi_{c}$.
-$\Phi$ is used for forward and backpropagation, where each gradient step applies the gradients calculated for $\Phi$ over $\Phi_{c}$.
-The latter is then binarized (without gradient tracking) to produce $\Phi$ for the next iteration.
-To demonstrate inherent robustness of parameter-space classifiers over voxel-grid data, in section 4.2.1 we first exemplify the efficacy of signal-domain attacks using BVA.
-Then, in section 4.2.2 we employ BVA in parameter-space attacks and demonstrate increased adversarial robustness.
-We emphasize that equation 4 is a specific instance of equation 2, and therefore all parameter-space attacks previously proposed in section 3 can be performed on voxel-grids using BVA.
+To perform the required binary optimization of $\delta$ we follow Iliadis et al.
+(2020), elaborated in section A.3.
 4 Results
-In the following sections we demonstrate the inherent robustness of Parameter-Space classifiers over 3 different datasets - MNIST ((Deng, 2012)), Fashion-MNIST ((Xiao et al., 2017)) and ModelNet10 ((Wu et al., 2015)).
-MNIST and Fashion-MNIST are chosen to represent the prevalent and heavily-studied modality of 2D images.
-While current adversarial robustness and computer-vision literature tends to focus on more challenging and diverse datasets (e.g.
-CIFAR100, ImageNet), applicability of these datasets in the parameter-space classification setting is currently limited.
-This is because state-of-the-art parameter-space classification algorithms proposed thus far have yet to achieve classification satisfying classification accuracies on such complex datasets ((Navon et al., 2023a, b; Chen & Wang, 2022)).
-ModelNet is used to demonstrate our findings over one of the modalities where usage of INRs is naturally useful, due to commonly large dimensionality.
-Following the approach from (Dupont et al., 2022), for each dataset $\mathcal{X}$ we first optimize a set of modulation vectors $\{v_{x}\forall x\in\mathcal{X}\}$.
+In the following sections we experiment with 3 different datasets - MNIST (Deng (2012)), Fashion-MNIST (Xiao et al. (2017)) and ModelNet10 (Wu et al. (2015)).
+Following the approach from Dupont et al.
+(2022), for each dataset $\mathcal{X}$ we first optimize a set of modulation vectors $\{v_{x}\forall x\in\mathcal{X}\}$.
 Then, we train a single MLP classifier over this modulation dataset.
@@
 Attack formulations and results for 2D data are presented in section 4.1, and for 3D data in section 4.2.
-In section 4.4 we further demonstrate the added computational costs induced by performing adversarial attacks over deep weight-space classifiers, posing added practical challenges for potential attackers.
+In section C we further demonstrate the added computational costs induced by performing adversarial attacks over deep weight-space classifiers, posing added practical challenges for potential attackers.
 Full training details are further specified in section A.1.
@@
 Results indicate 3 major trends - First, both MNIST and Fashion-MNIST are adversarially robust.
-For MNIST, while the TMO and BOTTOM attacks are the only two that managed to incur any above-marginal decrease in classification accuracy, it is still far below the accuracy decrease of around $60\%$ that is traditionally observed in similarly-bounded PGD attacks over signal-domain classifiers without any robust training ((Wong et al., 2020)).
-While higher-bounded attacks over the Fashion-MNIST dataset have been more successful, we claim their success is substantially lower than the effect of a similarly constrained attack over a signal-domain classifier.
+For MNIST, while the TMO and BOTTOM attacks are the only two that managed to incur any above-marginal decrease in classification accuracy, it is still far below the accuracy decrease of around $60\%$ that is traditionally observed in similarly-bounded PGD attacks over signal-domain classifiers without any robust training (Wong et al.
+(2020)).
+While higher-bounded attacks over the Fashion-MNIST dataset have been more successful, their success is substantially lower than the effect of a similarly constrained attack over a signal-domain classifier.
 To show this, we train four baseline signal-domain classifiers and attack with PGD using similar $\mathcal{L}_{\infty}$ norm constraints and PGD iteration number.
-We use Vision-Transformers (ViTs) (Dosovitskiy, 2020) due to their outstanding record in natural image classification (Han et al., 2022), and two convolutional models for their ubiquity and foundational role in computer-vision.
-To experiment with models of different sizes we train a ViT-Large model and a Resnet18, as well as a ViT-Base model and a compact,2-layer CNN.
-Full architecture details are included in Section A.3.2.
+For a representative pool of common signal-domain classifiers we use ViT-Base/16, ViT-Large/16 (Dosovitskiy (2020)), a Resnet18 and a smaller 2-layer CNN.
+Full architecture details are in section A.4.2.
 We compare the robust accuracy of these four Signal-space classifiers under standard 100-iteration PGD attack, with that of our parameter-space classifier subjected to all attacks from sections 3.3.1, 3.3.2, 3.3.3, 3.3.4 and 3.3.5.
-Results, shown in figure 3, show parameter-space classifiers show significantly higher robustness, especially for lower attack bounds.
-The second indicated trend is that we observe that the full PGD attack demonstrates lower adversarial prowess compared to BOTTOM and TMO attacks in all experiments.
+Results, shown in figure 2(a), indicate parameter-space classifiers show significantly higher robustness, especially for lower attack bounds.
+Furthermore, we observe that the full PGD attack demonstrates lower adversarial prowess compared to BOTTOM and TMO attacks in all experiments.
 We attribute this to the fact that in full-PGD the numerous steps of second-order optimization required for standard PGD in our case encourage vanishing gradients phenomena, namely attenuating gradient information backpropagated through the modulation optimization process $R$.
-These difficulties are added to the computational encumbrance of running full PGD (see section 4.4), and further stress the need in initial methods developed here (sections 3.3.2&3.3.3), as well as further related research.
 While computationally lightest, implicit differentiation proves less potent than most explicit methods.
 We attribute this to the strong dependence of this method in fulfillment of specified optimality conditions.
-Lastly, we conclude that the INR-domain attack (ICOP) failed to impose any actual effect on classification accuracy.
-We attribute this result to the heuristic nature of the approaches used to constrain the INR-domain perturbation’s response in the signal domain (section 3.3.4).
-We still include it within our suite of proposed attacks, as it proved more efficient when initiated over 3D data (section 4.2).
-Figure 3: Robust accuracy under $\mathcal{L}_{\infty}$ norm constraints - for parameter-space (solid curves) and signal-space (dotted curves) classifier attacks.
+(a) Fashion-MNIST robust accuracy in $\mathcal{L}_{\infty}$.
+(b) ModelNet10 robust accuracy in $\mathcal{L}_{0}$.
+Figure 2: Robust accuracy for parameter-space (solid curves) and signal-space (dotted curves) classifiers.
 Table 1: MNIST & Fashion-MNIST Clean/Robust Classification Accuracy - for any $\varepsilon$ value, $\mathcal{L}_{\infty}$ attack bound (equation 2) is $B=\frac{\varepsilon}{256}$.
@@
 Nonetheless, practical advantages of using INRs for 2D data are usually limited, since common grid representations suffice for most common tasks.
-Data modalities of larger dimensions, such as 3D data, however, more commonly gain from using INRs for achieving downstream tasks, including classification ((Jiang et al., 2020; Molaei et al., 2023; Niemeyer et al., 2020)).
-We use the ModelNet 10 dataset ((Wu et al., 2015)) consisting of 4900 3D CAD models given in Mesh-grid representation.
+Data modalities of larger dimensions, such as 3D data, however, more commonly gain from using INRs for achieving downstream tasks, including classification (Jiang et al.
+(2020); Molaei et al.
+(2023); Niemeyer et al.
+(2020)).
+In our experiments we use the ModelNet 10 dataset (Wu et al.
+(2015)) consisting of 4900 3D CAD models given in Mesh-grid representation.
 4.2.1 Signal Space Attack
 To demonstrate inherent adversarial robustness over 3D data under the BVA attack (section 3.4), we must first demonstrate its prowess in attenuating signal-domain classification accuracy.
-To do so we train a voxel-grid classification model based on Point-Voxel CNN layers proposed by (Liu et al., 2019b).
-We train the classifier over ModelNet10, converting every mesh-grid sample to a voxel-grid of dimension $15\times 15\times 15$ using Trimesh ((Dawson-Haggerty et al., )).
-Then, we attack this trained model with BVA for a wide range of $\mathcal{L}_{0}$ bounds and assess attacked classification accuracy.
-Results are displayed in figure 4, showing BVA is efficient over the tested signal-space classifier, inducing exponential decay of classification accuracy with growth in number of flipped voxel bits.
-Figure 4: BVA robust accuracy for signal-space classifier and weight-space classifier - under $\mathcal{L}_{0}$ norm constraints.
+To do so we train a voxel-grid classification model based on Point-Voxel CNN layers proposed by Liu et al.
+(2019b).
+We train the classifier over ModelNet10 (full optimization details brought in section A.1.2) and then attack this trained model with BVA for a wide range of $\mathcal{L}_{0}$ bounds and assess attacked classification accuracy.
+Results are displayed in figure 2(b), showing BVA is efficient over the tested signal-space classifier, inducing exponential decay of classification accuracy with growth in number of flipped voxel bits.
 4.2.2 Parameter Space Attack
-Figure 4 compares the robust accuracy of the signal domain (black) and parameter-space classifiers, where the parameter-space classifier is attacked with TMO, BOTTOM, ICOP, and implicit attacks.
-We note that we do not experiment with the full PGD attack in parameter-space for the 3D INR classifier, since to optimize INR for the ModelNet10 dataset with sufficient input signal fidelity, we had to perform 500 modulation optimization iterations per sample.
+Figure 2(b) compares the robust accuracy of the signal domain (black) and parameter-space classifiers, where the parameter-space classifier is attacked with TMO, BOTTOM, ICOP, and implicit attacks.
+We note that we do not experiment with the full PGD attack in parameter-space for the 3D INR classifier, since to optimize INR for the ModelNet10 dataset with sufficient input signal fidelity, we found it necessary to perform 500 modulation optimization iterations per sample.
 While computation times for doing so are not very high (several seconds), backpropagation through these 500 optimization steps calls for infeasible memory and compute time requirements, and also lead to severe vanishing/exploding gradient issues.
-We further elaborate on these considerations in Section 4.4.
+We further elaborate on these considerations in appendix C.
 While the parameter-space classifier exhibits lower clean accuracy (73% in parameter-space against 82% in signal space), results demonstrate conclusive increased adversarial robustness for the signal-space classifier.
@@
 The second is that for the attacked weight-space classifier, the decrease in accuracy attenuation across bound increase saturates significantly faster.
-The data from Figure 4 is presented numerically in Table 2.
+The data from fig.
+2(b) is presented numerically in table 2.
 4.3 Qualitative Results
 In this section we provide qualitative evidence for the inherent robustness of parameter-space classifiers.
-Adversarial attacks are known to effectively exploit the amplifications of small input-space perturbations (Goodfellow, 2014; Haleta et al., 2021).
 As previously discussed, we argue that the enhanced robustness of parameter-space classifiers arises from the attacker’s inability to directly perturb the classifier’s input.
-Instead, the attacker must optimize a constrained signal-domain adversarial perturbation for which the response received as input to the classifier (after modulation optimization Equation 1) is adversarially potent.
+Instead, the attacker must optimize a constrained signal-domain adversarial perturbation for which the response received as input to the classifier (after modulation optimization eq.
+1) is adversarially potent.
 This is deemed to be substantially more challenging for the attacker, as modulation optimization tends to attenuate adversarial prowess fed as input to the classifier.
-To illustrate this phenomenon, we report the mean amplification, measured as the difference in activation values between perturbed and clean input data, for each layer in our pipeline (Figure 2).
-These measurements, shown in Figure 5, were calculated using the MNIST dataset, with every sample propagated through the pipeline after being perturbed by each attack in our suite with a constraint of $\varepsilon=\frac{32}{256}$.
+To illustrate this phenomenon, we report the mean amplification, measured as the difference in activation values between perturbed and clean input data, for each layer in our pipeline (fig.
+1).
+These measurements, shown in fig.
+3, were calculated for classification over the MNIST dataset, with every sample propagated through the pipeline after being perturbed by each attack in our suite with a constraint of $\varepsilon=\frac{32}{256}$.
 To further validate the relationship between attack success and amplification, we differentiate between robust and non-robust samples.
@@
 This supports our claim that modulation optimization effectively obfuscates the adversarial potency of perturbations applied in the signal domain.
-Importantly, this effect is not observed for signal-domain classifiers, as we show in Figure 6.
+Importantly, this effect is not observed for signal-domain classifiers, as we show in fig.
+5.
 Results for other attack constraints and datasets are reported in appendix B.
 Figure 3: Adversarial amplification across pipeline layers - showing modulation optimization’s attenuation of adversarial amplification.
-In Figure 1 we demonstrate the affect of adversarial perturbations applied in signal-domain over the latent representations attained from perturbed data, and compare between parameter-space classifiers and traditional signal-space classifiers.
-To this end, as a signal-domain baseline, we first train a convolutional autoencoder over the MNIST dataset.
-For a fair comparison, this autoencoder is trained disjointly from the downstream classification task.
+In fig.
+4 we demonstrate the affect of adversarial perturbations applied in signal-domain over the latent representations attained from perturbed data, and compare between parameter-space classifiers and traditional signal-space classifiers.
+To this end, as a signal-domain baseline, we first train a convolutional autoencoder over the MNIST dataset (section A.4.2), disjointly from the downstream classification task.
 We then train a classifier over the learned latent space, and finally perform 100 iterations of signal-domain adversarial attack over this classifier.
-The bottom half of Figure 1 illustrates t-SNE projection of the trained autoencoder’s latent representations before (left) and after (right) the adversarial attack.
+The bottom half of fig.
+4 illustrates t-SNE projection of the trained autoencoder’s latent representations before (left) and after (right) the adversarial attack.
 The top half of the plot corresponds to parameter-space classifiers, showing t-SNE projections of INR modulations optimized for clean (right) data and for data attacked with our full-PGD attack (section 3.3.1.
-Prior to the attacks, the latent space of both signal and parameter-space domains sis well-structured, with clear separation between the different classes.
+Prior to the attacks, the latent space of both signal and parameter-space domains is well-structured, with clear separation between the different classes.
 However, after the attack, the model operating strictly in signal-domain (i.e.
@@
 Adversarial attacks optimized for the parameter-space classifier, on the other hand, seem inconsequential to the input-space structure of the parameter-space classifier, including class separability.
-These conclusions are well-correlated with those from Figure 5.
-4.4 Computational Resource Analysis
-In sections 4.1,4.2 we provide evidence for the adversarial robustness of parameter-space classifiers based on classification accuracies under the presence of adversarial perturbations.
-In this section, we further support these claims from a different, practical perspective.
-We show that orthogonally to the efficacy of adversarial attacks over parameter-space classifiers, attacking these classifiers also poses significant added computational difficulties for the potential attacker.
-These difficulties stem from the optimization process inherent within the classification pipeline’s inference process.
-To illustrate this, we measure the computation times required for a single modulation optimization step during a single PGD iteration on an NVIDIA RTX-2080 GPU.
-Our results show that for the MNIST dataset with 512-dimensional modulation vectors, the compute time per step is approximately 0.015 seconds, while for the ModelNet10 dataset with 2048-dimensional modulation vectors, it is 0.02 seconds.
-Even for a small number of PGD steps, increasing the number of modulation optimization steps – an aspect controlled by the model, not the attacker – increases attack optimization times within several orders of magnitude.
-Importantly, this increase does not significantly impact the computation time required for ”clean” (unperturbed) inference.
-For instance, with 100 modulation optimization steps set by the model to solve Equation 1, ”clean” inference would require 1.5 seconds per sample, while perturbation optimization of 100 PGD iterations would require the attacker to invest 150 seconds per-sample.
-On top of computation time increase, attacking parameter-space classifiers via explicit differentiation also poses a challenge for the attacker in terms of memory consumption.
-Tested on the same hardware, we find that GPU memory consumption for the full PGD attack over the ModelNet10 classifier grows linearly with number of modulation optimization steps, with a slope of $0.1705[\frac{GB}{Mod.Steps}]$.
-The corresponding coefficient for the smaller MNIST data is $0.049[\frac{GB}{Mod.Steps}]$.
-Namely, in order to perform a full PGD (Section 3.3.1) attack over the ModelNet10 data, every additional modulation optimization step included in ”clean” model training induces roughly 170 additional megabytes for the attacker’s backpropagation.
-Backpropagating through the 500 modulation optimization steps used in creation of the ModelNet10 INR dataset in our experiments would require 85 Gigabytes of GPU memory, greatly increasing optimization costs.
-The increase in modulation optimization steps holds no more than marginal prices for the ”innocent” (non-attacker) user in INR dataset optimization, as this process does not require 2nd-order differentiation.
-Implicit differentiation (Section 3.3.5) offers the attacker constant-memory attacks (in our experiments - 526 MB for MNIST, 986 MB for ModelNet10), however at the cost of higher sensitivity to optimality condition satisfaction (as shown in Section 4.1).
-5 Discussion
-5.1 Conclusion
-In this work, we’ve explored the adversarial robustness of parameter-space classifiers.
-We’ve presented the challenges involved in translating common PGD-based attacks onto models performing classification directly in deep parameter-space, and addressed these challenges by developing a novel set of five different adversarial attacks designed for parameter-space classification, applying perturbations both in the signal domain and in parameter space.
-We then tested these attack over three different datasets from two popular modalities, and shown that parameter-space classifiers exhibit an increased inherent adversarial robustness, compared to their counterparts classifying in signal-domain.
-We’ve further presented the notion of ”robustness” stemming from the potentially large added computational resources required for certain cases of parameter-space classifier attacks.
-5.2 Limitations and Future Work
-While we consider our work an important initial study of the currently under-explored topic of adversarial robustness in parameter-space classification, several notable limitations remain.
-First, in this work we consider a relatively limited set of simple datasets from both 2D and 3D modalities.
-This choice mainly stems from the fact that current state-of-the-art parameter space classifiers have yet to achieve adequate ”clean” (non-attacked) classification results over more challenging datasets ((Navon et al., 2023a, b)).
-Since the focus of our work is adversarial robustness of parameter-space classifiers, experimenting with our proposed attacks over classifiers with low ”clean” accuracy would make assessment of our work less reliable.
-We hope future exploration of methods for better non-adversarial parameter-space classification would expand our set of potential datasets for robustness exploration.
-Second, we note that the adversarial robustness potentially attained when using parameter-space classifiers often comes with the cost of reduced classification accuracy, as shown in section 4.
-Finally, our work focuses on inherent adversarial robustness, and does not discuss active defenses.
-We deem our findings about inherent robustness an encouraging starting point for future research in development of adversarial defenses and robust modeling.
+These conclusions are well-correlated with those from fig.
+3.
+Figure 4: t-SNE projection of modulation vectors fitted for clean and adversarially-perturbed data - for parameter-space and signal-space classifiers.
+5 Conclusion
+In this work, we investigate the adversarial robustness of parameter-space classifiers.
+We highlight the challenges in adapting commonly used PGD-based attacks to models that perform classification directly in the deep parameter space.
+To address these challenges, we propose a novel set of five distinct adversarial attacks tailored for parameter-space classification.
+Our results demonstrate that parameter-space classifiers exhibit enhanced inherent adversarial robustness compared to those operating strictly in the signal domain.
+While this study serves as a significant initial exploration of adversarial robustness in parameter-space classification—a relatively under-explored area—we acknowledge several limitations.
+First, current state-of-the-art parameter-space classifiers still struggle to achieve satisfactory ”clean” (non-attacked) classification performance on more complex datasets (Navon et al.
+(2023a; b)).
+Consequently, we focus on a limited set of relatively simple datasets from both 2D and 3D modalities, where ”clean” accuracy is sufficient for a reliable evaluation of adversarial attack effects.
+We hope that future advancements in non-adversarial parameter-space classification will broaden the range of datasets available for robustness studies.
+Second, we note that the adversarial robustness observed in parameter-space classifiers often comes at the expense of reduced classification accuracy, as demonstrated in section 4.
 Appendix A Training Regime, Hyperparameters & Technical Details
@@
 All experiments described in this section have been conducted on a single NVIDIA RTX-2080 GPU.
-Following (Dupont et al., 2022), for each of the three datasets considered in this paper, we first train a SIREN (Sitzmann et al., 2020) jointly over the entire training set.
+Following Dupont et al.
+(2022), for each of the three datasets considered in this paper, we first train a SIREN Sitzmann et al.
+(2020) jointly over the entire training set.
 For MNIST and Fashion-MNIST we optimize modulation vectors of dimension $512$ for 6 epochs of 3 modulation optimization steps per-sample, with a learning-rate of $0.000005$ and a batch size of $128$.
@@
 A.1.2 ModelNet10
-For ModelNet10 we similarly follow (Dupont et al., 2022), however we first convert the Mesh-grid data onto a voxelgrid, as previously mentioned (Section 3.4).
+For ModelNet10 we similarly follow Dupont et al.
+(2022), however we first convert the Mesh-grid data onto a voxelgrid, as previously mentioned (section 3.4).
 Each voxelgrid is resampled to dimension $15\times 15\times 15$.
@@
 We use the Adam optimizer in all experiments.
-A.3 Baseline Models
-A.3.1 Classifier Baselines
+A.3 Binary Optimization in BVA
+In BVA eq. 5 we formulate the attacker objective as:
+$\max_{\delta}\quad$ $\|M_{\psi}(x\oplus\delta)\|\text{s.t.}\|p\|_{0}=B$ (5)
+Since voxel-grid data is represented by binary mask, our optimized perturbation $\delta$ in this case is a binary mask governing the bit-flip imposed over the original, ”clean” voxel-grid data.
+This formulation calls for the incorporation of binary optimization of the mask $\delta$.
+Trivial binarization of $\delta$ (e.g.
+by thresholding) is non-differentiable.
+To optimize $\delta$ as a binary mask we follow the binary optimization technique presented in Iliadis et al.
+(2020).
+This method maintains two congruent masks - a binary mask $\Phi$ and a real mask $\Phi_{c}$.
+$\Phi$ is used for forward and backpropagation, where each gradient step applies the gradients calculated for $\Phi$ over $\Phi_{c}$.
+The latter is then binarized (without gradient tracking) to produce $\Phi$ for the next iteration.
+In our experiments, we optimize under varying $\mathcal{L}_{0}$ constraints (fig. 2(b)), using a learning-rate of $0.01$.
+A.4 Baseline Models
+A.4.1 Classifier Baselines
 In this section we specify implementation and training details related to the signal-space classifier baseline CNN and ViT architectures used in section 4.
 In section 4.1 we train 4 signal-domain classifiers – a small CNN, a ResNet18, ViT-Large/16 and Vit-Base/16.
-For the ViTs we leverage pre-trained models from (Wightman, 2019), and finetune for our classification task using a a linear probe prior to a applying Softmax.
+For the ViTs we leverage pre-trained models from Wightman (2019), and finetune for our classification task using a a linear probe prior to a applying Softmax.
 For the Resnet we train from scratch, and project encoded data with a linear probe followed by a Softmax.
@@
 For all baselines we train over the Fashion-MNIST training set for 5 epochs with learning-rate of 0.0001.
-A.3.2 Autoencoder Baselines
+A.4.2 Autoencoder Baselines
 In section 4.3, we present a qualitative analysis of adversarial perturbation amplification across pipeline layers, along with t-SNE projections of modulation spaces before and after adversarial perturbation.
@@
 After the autoencoder is trained, we freeze it and train a linear classifier with 2 hidden layers and a single ReLU activation on top of it.
-The frozen encoder layers acts as the INR counterpart in Figure 5, and the MLP head is parallelized to the parameter-space classifier.
+The frozen encoder layers acts as the INR counterpart in fig.
+3, and the MLP head is parallelized to the parameter-space classifier.
 We use a latent dimension of $128$ for all encoders.
@@
 B.1.1 MNIST
-Figure 6 illustrates the mean amplification difference measured through model activation layers over MNIST data, similarly to Figure 5, however for the signal-domain classifier trained according to section A.3.2.
+fig.
+5 illustrates the mean amplification difference measured through model activation layers over MNIST data, similarly to fig.
+3, however for the signal-domain classifier trained according to section A.4.2.
 This model uses a 3-layer convolutional encoder, followed by a classifier with a single activation.
 Clean classifier accuracy is $98.8\%$, robust accuracy is $11.28\%$.
-Differently from similar results displayed for parameter-space classifiers in Figure 5, upon transition from encoder to classifier layers (3rd layer, marked by the black vertical line), the adversarial amplification is not attenuated, but rather drastically inflated.
+Differently from similar results displayed for parameter-space classifiers in fig.
+3, upon transition from encoder to classifier layers (3rd layer, marked by the black vertical line), the adversarial amplification is not attenuated, but rather drastically inflated.
 Figure 5: Signal-Domain Amplification - through encoder and downstream classifier layers
-Figure 7 showcases adversarial amplification (similarly to Figure 5) across different attack bound for the full-PGD attack. Adversarial obfuscation drawn from Figure 5 arises here as well.
+fig. 6 showcases adversarial amplification (similarly to fig. 3) across different attack bound for the full-PGD attack. Adversarial obfuscation drawn from fig. 3 arises here as well.
 Figure 6: Signal-Domain Amplification - through encoder and downstream classifier layers, across different constraints $\frac{\varepsilon}{256}$ for full PGD attack
 B.1.2 Fashion-MNIST
-Figure 8 illustrates adversarial amplification across pipeline layers, similarly to Figure 5.
+fig.
+7 illustrates adversarial amplification across pipeline layers, similarly to fig.
+3.
 Trends of correlation between amplification and adversarial success, as well as evident obfuscation of adversarial affect between INR and classifier segments, are consistent with findings from section 4.3.
 Figure 7: Adversarial amplification across pipeline layers - showing modulation optimization’s attenuation of adversarial amplification.
-Similarly to Figure 6, we test amplification patterns of a signal-domain classifier (Section A.3.2) for the Fashion-MNIST dataset as well.
+Similarly to fig.
+5, we test amplification patterns of a signal-domain classifier (section A.4.2) for the Fashion-MNIST dataset as well.
 We again observe consistency across both datasets - the signal domain classifiers proves to be much more sensitive to adversarial examples propagated through the encoder-classifier pipeline.
 Figure 8: Signal-Domain Amplification - through encoder and downstream classifier layers
-Figure 10 exhibits the affect of adversarial attacks in signal-domain over the underlying modulation vectors (for parameter-space classifiers) and latent vectors (for signal-space classifiers) received as classifier input.
-Similarly to Figure 1, for both parameter-space and signal-space model, a well-formed latent space is observed before the incorporation of adversarial perturbation.
+fig.
+9 exhibits the affect of adversarial attacks in signal-domain over the underlying modulation vectors (for parameter-space classifiers) and latent vectors (for signal-space classifiers) received as classifier input.
+Similarly to fig.
+4, for both parameter-space and signal-space model, a well-formed latent space is observed before the incorporation of adversarial perturbation.
 After attacking with PGD, the corresponding modulation vectors fed into the parameter-space classifier remain almost unchanged, whilst the latent space observed by the signal-domain classifier is heavily deformed.
@@
 In this section we report additional results for the ModelNet10 dataset.
-In Table 2 we report numerical results for the data from Figure 4 The first two columns represent the same attack constraints in absolute bit numbers ($\varepsilon$) and bit numbers relative to the entire volume (matching horizontal axis in figure 4).
+In table 2 we report numerical results for the data from fig.
+2(b) The first two columns represent the same attack constraints in absolute bit numbers ($\varepsilon$) and bit numbers relative to the entire volume (matching horizontal axis in figure 2(b)).
 Table 2: ModelNet10 Clean/Robust Classification Accuracy - for any given $\varepsilon$ value, attack bound from equation 2 is $B=\varepsilon$ in $\mathcal{L}_{0}$ (i.e. number of flipped signal bits).
@@
 200 0.06 0.734 0.594 0.576 0.41 0.613
+Appendix C Computational Resource Analysis
+In sections 4.1,4.2 we provide evidence for the adversarial robustness of parameter-space classifiers based on classification accuracies under the presence of adversarial perturbations.
+In this section, we further support these claims from a different, practical perspective.
+We show that orthogonally to the efficacy of adversarial attacks over parameter-space classifiers, attacking these classifiers also poses significant added computational difficulties for the potential attacker.
+These difficulties stem from the optimization process inherent within the classification pipeline’s inference process.
+To illustrate this, we measure the computation times required for a single modulation optimization step during a single PGD iteration on an NVIDIA RTX-2080 GPU.
+Our results show that for the MNIST dataset with 512-dimensional modulation vectors, the compute time per step is approximately 0.015 seconds, while for the ModelNet10 dataset with 2048-dimensional modulation vectors, it is 0.02 seconds.
+Even for a small number of PGD steps, increasing the number of modulation optimization steps – an aspect controlled by the model, not the attacker – increases attack optimization times within several orders of magnitude.
+Importantly, this increase does not significantly impact the computation time required for ”clean” (unperturbed) inference.
+For instance, with 100 modulation optimization steps set by the model to solve eq.
+1, ”clean” inference would require 1.5 seconds per sample, while perturbation optimization of 100 PGD iterations would require the attacker to invest 150 seconds per-sample.
+On top of computation time increase, attacking parameter-space classifiers via explicit differentiation also poses a challenge for the attacker in terms of memory consumption.
+Tested on the same hardware, we find that GPU memory consumption for the full PGD attack over the ModelNet10 classifier grows linearly with number of modulation optimization steps, with a slope of $0.1705[\frac{GB}{Mod.Steps}]$.
+The corresponding coefficient for the smaller MNIST data is $0.049[\frac{GB}{Mod.Steps}]$.
+Namely, in order to perform a full PGD (section 3.3.1) attack over the ModelNet10 data, every additional modulation optimization step included in ”clean” model training induces roughly 170 additional megabytes for the attacker’s backpropagation.
+Backpropagating through the 500 modulation optimization steps used in creation of the ModelNet10 INR dataset in our experiments would require 85 Gigabytes of GPU memory, greatly increasing optimization costs.
+The increase in modulation optimization steps holds no more than marginal prices for the ”innocent” (non-attacker) user in INR dataset optimization, as this process does not require 2nd-order differentiation.
+Implicit differentiation (section 3.3.5) offers the attacker constant-memory attacks (in our experiments - 526 MB for MNIST, 986 MB for ModelNet10), however at the cost of higher sensitivity to optimality condition satisfaction (as shown in section 4.1).